OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix?

OpenStudy (anonymous):

if the ten pound cost $3.90, how much is the total cost?

OpenStudy (anonymous):

if you put the number of pounds of cashews as \(x\) then they cost \(5.6x\) for the cashews. the total is 10 pounds, so the number of pounds of peanuts must be \(10-x\) at a cost of \(2.30(10-x)\) for a total cost of \(5.60x+2.30(10-x)\) and since the total cost is \(10\times 3.60=39\) set \[5.6x+2.3(10-x)=39\] and solve for \(x\)

OpenStudy (anonymous):

easier to solve if you remove the annoying decimal by multiplying both sides by 10 and solve \[56x+23(10-x)=390\] you good from there ?

OpenStudy (anonymous):

whoa i am wrong sorry hold the phone

OpenStudy (anonymous):

okay thasnk you.. so one more question, 1/6, 1/3, 1/2, 2/3 would the common difference be 1/6? & okay, thank you

OpenStudy (anonymous):

it is $3.29 not $3.9 so you need to solve \[56x+23(10-x)=329\]

OpenStudy (anonymous):

to your second question, yes, the common difference is \(\frac{1}{6}\)

OpenStudy (anonymous):

okay, thank yoyu

OpenStudy (anonymous):

could you aklso set it up as 3.29x + 5.60x = 10 /